OpenStudy (anonymous):

can someone please help me solve this? thank you! 5(x+2)-5x=40

OpenStudy (anonymous):

use: A(x+B)=Ax+AB,But constant on side ,but variable (x) on other side and solve for x.

OpenStudy (anonymous):

Are you sure you wrote the problem correctly? When you simplify this you get: \[5(x+2) - 5x = 40\] \[5x + 10 - 5x = 40\] \[10 \neq 40\]

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

and yes thats what i have

OpenStudy (anonymous):

no solution